OpenStudy (anonymous):

Suppose a country's population in 1980 was 210 million. In 1990 it was 225 million. Using the exponential growth formula, P=Ae^kt, estimate the country's population in 2000

OpenStudy (anonymous):

P=Ae^kt 225 = 210 * e^(k*(1990-1980) 225/210=e^10k ln(225/210)=10k k=ln(225/210)/10=0.0069 P = 210*e^(0.0069t) for 2000 ===> t = 2000-1980=20 P = 210*e^(0.0069*20) P=241.0749=241